I need help with question D, basically interpreting the question and it’s answer. What units are avg labour productivity measured in ? And what does it imply ?
Ignore the 2nd table ( was for another question )
2017 9018 100 200 Output (TVs) Employment (Computers) Employment (TVs) Price per Computer 1000 500 $1,000 $900 101 1100 400 $400 $500 Table 1: Question 2 Value GDP NFP (Net Factor Payments) C (Consumption) G (Government Expenditures) PDI (Personal Disposable Income) I (Investment) T (Tares) 2000 100 200 160 1400 1600 640 TR (Government Transfers)130 Table 2: Question 3 Question 2: Employment, Output, Productivity and Inflation (20 Marks) Assume that in the Economy only two goods are produced: Computers and TVs. Macroeconomic information for this economy is given in Table 1 a. How much is Nominal GDP in 2017 and 2018? What is the percentage change? percentage change? the base year to 2018? 20187 b. How much is Real GDP in 2017 and 2018, by considering 2017 as the base year? What is the c. How much is the GDP deflator in the two years? By what percentage does the price level change from d. What was the growth rate of average labour productivity for the whole economy between 2017 and
Answer #1

Hi

Here, the labour productivity would be measured as Output per Worker.
Hence, in 2017- Output (total) = 100 + 200 = 300, and employment (total) = 1000 + 500 = 1500.
Hence, Output per worker for 2017 =300/1500 = 0.2

Now, in 2018- Output (total) = 101 + 201 = 302, and employment (total) = 1100 + 400= 1500.
Hence, Output per worker for 2018 =302/1500 = 0.201

Hence, labour productivity has increased by 0.001 = 0.1%